A company is expected to increase dividends by 20% in year 1, and by 15% in year 2, after that, dividends will increase at a rate of 5% per year indefinitely. If the last dividend was $1, and the required return is 15%, what is the current price of the stock? (Hint: current price will be from NPV)

a. $6.90

b. $8.67

c. $9.66

d. $11.10

e. $13.04
Answer #1

D1=(1*1.2)=1.2

D2=(1.2*1.15)=1.38

Value after year 2=(D2*Growth rate)/(Required return-Growth rate)

=(1.38*1.05)/(0.15-0.05)=$14.49

Hence current price=Future dividends*Present value of discounting factor(15%,time period)

=1.2/1.15+1.38/1.15^2+14.49/1.15^2

which is equal to

=$13.04(Approx).